# CLiMB: A Domain-Informed Novelty Detection Clustering Framework for Scientific Discovery

### v1

@@ 4.2 Data Preparation and Preprocessing
Energy Decorrelation:
Systematic biases in energy estimation can introduce artificial correlations. To mitigate this, we applied a statistical shuffling technique. We generated 1,000 permutations of the energy feature vector $\mathbf{e}$; the mean of these shuffled vectors was then subtracted from the original energy values:
$\mathbf{e}^{\prime}=\mathbf{e}-\mathbb{E}[\text{shuffle}(\mathbf{e})]_{1000}$
This procedure effectively removes the mean background energy level while preserving local kinematic structures.
@@ 6 Conclusions
Acknowledgments and Disclosure of Funding
Support to this study has been provided by INAF Mini-Grant (PI: Tatiana Muraveva), by the Agenzia Spaziale Italiana (ASI) through contract ASI 2018-24-HH.0, its Addendum 2018-24-HH.1-2022 and contract ASI 2025-10-H.00, and by Premiale 2015, MIning The Cosmos - Big Data and Innovative Italian Technology for Frontiers Astrophysics and Cosmology (MITiC; P.I.B.Garilli). This work uses data from the European Space Agency mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. More information about this disclosure can be found on the JMLR website.
@@ Appendix A Hyperparameter Optimization
For our proposed algorithm, CLiMB, we employed a hierarchical, two-stage optimization strategy tailored to its architecture:
@@
This hierarchical approach ensures that each phase of CLiMB is fine-tuned for its specific task. The final, optimal hyperparameters used for the benchmark analysis presented in this paper are detailed in Table 2. The complete code for the optimization routines, along with the scripts to reproduce all results, is publicly available in our GitHub repository [^6] https://github.com/LorenzoMonti/RRLs-Clustering-Benchmark.

### v2

@@ 4.2 Preprocessing and Feature Engineering
Energy Centering:
To mitigate biases related to the absolute scale of total energy and to facilitate the computation of the Mahalanobis distance, we centered the energy distribution. Specifically, we subtracted the mean background energy level from the original values:
$\mathbf{e}^{\prime}=\mathbf{e}-\mu_{e}$
This procedure ensures that the feature space is centered around the origin regarding the energy component.
@@ 5.1 Knowledge Recovery and Baseline Comparison
This performance gap highlights the critical role of CLiMB’s radial constraint. Unlike C-DBSCAN, which allows cluster centers to migrate freely based on local density bridges, often leading to ’semantic drift’ where distinct structures merge via background noise, CLiMB utilises the seed centroids as physical anchors. By constraining the centroid displacement ($\tau_{r}$), the algorithm is forced to adapt the cluster shape (covariance) to the local data distribution without losing the kinematic identity ($E,L_{z}$) defined by the prior literature. This ensures that the known structures are tightly recovered, leaving the residual space clean for the subsequent exploratory phase.
@@ 6 Conclusions
Acknowledgements
Support to this study has been provided by INAF Mini-Grant (PI: Tatiana Muraveva), by the Agenzia Spaziale Italiana (ASI) through contract ASI 2018-24-HH.0, its Addendum 2018-24-HH.1-2022 and contract ASI 2025-10-H.00, and by Premiale 2015, MIning The Cosmos - Big Data and Innovative Italian Technology for Frontiers Astrophysics and Cosmology (MITiC; P.I.B.Garilli). This work uses data from the European Space Agency mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
@@ Appendix A Hyperparameter Optimization
For our proposed algorithm, CLiMB, we employed the Mahalanobis distance metric for the constrained phase to naturally model the ellipsoidal correlations typical of kinematic streams ($E$-$L_{z}$). We then used a hierarchical, two-stage optimization strategy:
@@
This hierarchical approach ensures that each phase of CLiMB is fine-tuned for its specific task. The final, optimal hyperparameters used for the benchmark analysis presented in this paper are detailed in Table 2. The complete code for the optimization routines, along with the scripts to reproduce all results, is publicly available in our GitHub repository666github.com/LorenzoMonti/RRLs-Clustering-Benchmark.